OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)? y + 6 = 1(x + 3) y + 1 = −1(x + 2) y − 6 = −1(x + 3) y − 6 = 1(x − 3)

OpenStudy (anonymous):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-1 \right)\]